# A Unified Charge-Dependent Modulation Model for AMS-02 Proton and Antiproton Fluxes during Solar Minimum

### v1

@@ Abstract
We develop a unified charge-dependent solar modulation model by solving the three-dimensional Parker transport equation, incorporating a realistic wavy heliospheric current sheet to treat drift effects self-consistently. Using a local interstellar spectrum from GALPROP constrained by Voyager data, we fit the model to time-resolved proton and antiproton fluxes measured by the Alpha Magnetic Spectrometer - 02 (AMS-02) during the solar-quiet period (May 2011 – June 2022). To enable rapid parameter scans, we employ neural-network-based surrogate models to compute propagation and modulation matrices efficiently. The results demonstrate that the model simultaneously describes the observed proton and antiproton fluxes with physically reasonable parameters, providing a unified account of charge-dependent modulation.
@@ I Introduction
Galactic cosmic rays (GCRs) detected at the top of the atmosphere (TOA) originate from sources such as supernova remnants (SNRs). After propagating through the galaxy, they undergo modulation by the heliospheric magnetic field (HMF) upon entering the solar system. The injection, propagation and modulation processes collectively shape the GCR spectrum observed at TOA. Consequently, the interdependence of these processes complicates the interpretation of cosmic ray flux, especially in the GeV energy range, obscuring potential anomalous signals, such as secondary particles from dark matter with masses in the tens of GeV range linInvestigatingDarkMatter2019 ; cuiPossibleDarkMatter2017 ; yuanPropagationCosmicRays2017 ; zhuExplainingGeVAntiproton2022 .
Two strategies effectively disentangle the modulation process from the injection and propagation processes: examining the local interstellar spectrum (LIS) and studying temporal variations in the cosmic-ray (CR) spectrum. The CR LIS, entirely free of heliospheric modulation, solely reflects injection and propagation effects. However, only the two Voyager spacecrafts have traversed beyond the heliosphere to date cummingsGALACTICCOSMICRAYS2016 ; stoneCosmicRayMeasurements2019 . Their CR observations, spanning MeV to sub-GeV energies, remain the only available LIS measurements. Alternatively, temporal variations in the TOA spectrum can separate modulation effects, as the HMF varies over the 11-year solar cycle while injection and propagation processes remain constant. The AMS-02 and the Payload for Antimatter Matter Exploration and Light-nuclei Astrophysics (PAMELA) experiments have provided precise, time-resolved CR flux measurements over a full solar cycle adrianiTimeDependenceProton2013 ; amscollaborationTemporalStructuresElectron2023 ; amscollaborationTemporalStructuresPositron2023 ; aguilarPeriodicitiesDailyProton2021 ; amscollaborationAntiprotonsElementaryParticles2025 . Combined with Voyager’s sub-GeV LIS data, these measurements can be employed to illuminate modulation effects.
@@
Using recently published time-resolved fluxes of protons, antiprotons, electrons, and positrons measured by the AMS-02 collaboration from May 2011 to June 2022, we investigate drift effects more realistically amscollaborationAntiprotonsElementaryParticles2025 . We numerically solve the three-dimensional Parker transport equation for CR diffusion in the HMF via a time-backward stochastic method jiangNewScenarioSolar2023 . This approach, consistent with the SolarProp software kapplSOLARPROPChargesignDependent2016 , also incorporates a realistic wavy heliospheric current sheet (HCS) structure to model HCS drift, a component distinct from curvature and gradient drifts in the HMF peiThreedimensionalWavyHeliospheric2012 . By applying the LIS derived from the GALPROP package moskalenkoProductionPropagationCosmicRay1998 ; strongPropagationCosmicRayNucleons1998a , we demonstrate that this model provides a unified description of modulation effects for all the four elementary particle species, regardless of charge sign or mass, and is fitted to Voyager and AMS-02 measurements. Such a unified model can be further applied to study potential anomalous signals that are obscured in the GeV energy range of CR spectra.
@@ III.2 Diffusion Model
The three-dimensional Parker’s equation, with the complex structure described above, is solved using a backward-in-time stochastic differential equation (SDE) method. In this approach, pseudo-particles are traced backward from the Earth to the heliopause. The evolution of a particle’s position is governed by the SDE yamadaStochasticViewSolar1998 ; zhangMarkovStochasticProcess1999 ; koppStochasticDifferentialEquation2012 ; kapplSOLARPROPChargesignDependent2016 :
@@ IV.1 Machine Learning Model
Figure 2: Deep learning architecture for spectral data processing in space physics, featuring propagation matrix (PropMat) generation and analysis across spatial domains (Galaxy and Heliosphere) using an encoder-decoder framework.
@@
To handle multiple particle types, a subspace-based decoding scheme is implemented. The encoded latent vector is projected by a dedicated MLP into a particle-specific subspace of dimension $n_{\mathrm{sub}}$. The resulting vector serves as a set of coefficients, which are then fed into a coefficient-based matrix synthesis layer. In this layer, the coefficients linearly combine a set of learnable basis matrices to generate the final transformation matrix for the target particle type.
In this work, the subspace dimension $n_{\mathrm{sub}}$ is fixed at 32. This limited dimensionality constrains the matrix representation to a compact basis, encouraging the model to learn the principal features of the transformation. Such an approach is analogous to the Principal Component Analysis (PCA) algorithm jolliffePrincipalComponentAnalysis2014 , but substitutes its data-derived orthogonal basis with a set of learnable non-orthogonal basis matrices optimized for a custom objective. For physical transformation matrices, these principal features, typically limited in number, often correspond to low-frequency components. Consequently, the method inherently promotes smoothness in the learned transformations, ensuring a smooth output spectrum and thus supporting better generalization of the propagation result.
@@
$\text{Loss}_{\text{G}}=\sum_{j}\Bigl(\log\frac{\sum_{i}M_{\text{pred}}^{ij}\,\psi_{\mathrm{I}}^{i}}{\psi_{\mathrm{O}}^{j}}\Bigr)^{2}+\sum_{ij}\beta\log\bigl(M_{\text{dev}}^{ij}\bigr)^{2},$ (17)
where the MSE is evaluated on a logarithmic scale because the spectral values span several orders of magnitude. The hyperparameter $\beta$, which controls the regularization strength, is assigned with a relatively large value at first and decreased gradually.
@@
where, as in Equation (17), the MSE is also evaluated on a logarithmic scale. Since the typical energy change during heliospheric modulation is on the order of the modulation potential (approximately 1GeV1\ $\mathrm${GeV} for $|Z|=1$ particles), the energy-bin width in the transformation matrix must be considerably smaller than this scale. We therefore adopt a linear binning scheme with widths of 0.1GeV0.1\ $\mathrm${GeV} or 0.05GeV0.05\ $\mathrm${GeV}.
@@ V Results
Figure 6: Best-fit proton (left) and antiproton (right) fluxes. The gray line represents the LIS, while the colored solid lines show the modulated spectra for the corresponding Bartels rotations. Colored dashed lines indicate the corresponding modulated spectra from the FFA, with the modulation potential $\phi$ given in the legend.
@@
Figure 6 shows the best-fit local interstellar spectra for protons and antiprotons, along with the modulated spectra for several representative Bartels rotations, compared with Voyager and AMS-02 data. For comparison, the modulation result obtained using the FFA is plotted with dashed lines. The proton flux during Bartels rotation 2531 closely matches that during rotation 2543, whereas the antiproton flux is significantly higher in the latter. This difference results from the smaller tilt angle $\alpha$ in rotation 2543. During the studied period under positive magnetic polarity, as mentioned in Section III, protons reach Earth primarily through the polar region and are thus insensitive to changes in the tilt angle. In contrast, antiprotons detected at Earth are transported mainly via HCS drift. A larger tilt angle extends the drift path, enhances particle losses during transport, and consequently lowers the observed flux. Compared to the FFA, which requires distinct modulation potentials for the two Bartels rotations to account for such differences, our model incorporates tilt angle information and successfully reproduces the observed variations for both protons and antiprotons.

### v2

@@ Abstract
We develop a unified charge-dependent solar modulation model by solving the three-dimensional Parker transport equation, incorporating a realistic wavy heliospheric current sheet to treat drift effects self-consistently. Using a local interstellar spectrum from GALPROP constrained by Voyager data, we fit the model to time-resolved proton and antiproton fluxes measured by the Alpha Magnetic Spectrometer-02 (AMS-02) during the solar-quiet period (May 2011 to June 2022). To enable rapid parameter scans, we employ neural-network-based surrogate models to compute propagation and modulation matrices efficiently. The results demonstrate that the model simultaneously describes the observed proton and antiproton fluxes with physically reasonable parameters, providing a unified account of charge-dependent modulation.
@@ I Introduction
Galactic cosmic rays (GCRs) detected at the top of the atmosphere (TOA) originate from sources such as supernova remnants (SNRs). After propagating through the galaxy, they undergo modulation by the heliospheric magnetic field (HMF) upon entering the solar system. The injection, propagation and modulation processes collectively shape the GCR spectrum observed at TOA. Consequently, the interdependence of these processes complicates the interpretation of cosmic ray flux, especially in the GeV energy range, obscuring potential anomalous signals, such as the low-energy secondary particles from dark matter linInvestigatingDarkMatter2019 ; cuiPossibleDarkMatter2017 ; yuanPropagationCosmicRays2017 ; zhuExplainingGeVAntiproton2022 .
Two strategies effectively disentangle the modulation process from the injection and propagation processes: examining the local interstellar spectrum (LIS) and studying temporal variations in the cosmic-ray (CR) spectrum. The CR LIS, entirely free of heliospheric modulation, solely reflects injection and propagation effects. However, only the two Voyager spacecrafts have traversed beyond the heliosphere to date cummingsGALACTICCOSMICRAYS2016 ; stoneCosmicRayMeasurements2019 . Their CR observations, spanning MeV to sub-GeV energies, remain the only available LIS measurements. Alternatively, temporal variations in the TOA spectrum can separate modulation effects, as the HMF varies over the 11-year solar cycle while injection and propagation processes remain constant. The AMS-02, Payload for Antimatter Matter Exploration and Light-nuclei Astrophysics (PAMELA), and High Energy Cosmic-radiation Detection (HERD) experiments have provided precise, time-resolved CR flux measurements over a full solar cycle adrianiTimeDependenceProton2013 ; amscollaborationTemporalStructuresElectron2023 ; amscollaborationTemporalStructuresPositron2023 ; aguilarPeriodicitiesDailyProton2021 ; amscollaborationAntiprotonsElementaryParticles2025 ; xuPotentialConstrainingPropagation2023 . Combined with Voyager’s sub-GeV LIS data, these measurements can be employed to illuminate modulation effects.
@@
Beyond the phenomenological FFA model, more sophisticated models based on the time-backward stochastic differential equation (SDE) solution of Parker’s transport equation have been developed to simultaneously describe modulation effects on particles of different charge signs. These models, whether in two dimensions kapplSOLARPROPChargesignDependent2016 ; jiangNewScenarioSolar2023 or three potgieterModulationGalacticProtons2014 ; songNumericalStudySolar2021 ; duanScrutinizingImpactSolar2025 , can simulate both positively and negatively charged particles under identical physical conditions. Using the time-resolved fluxes of protons, antiprotons, electrons, and positrons measured by the AMS-02 collaboration from May 2011 to June 2022amscollaborationAntiprotonsElementaryParticles2025 , we investigate drift effects with such models, and adopt a recent 3D heliospheric current sheet (HCS) drift model derived from a realistic wavy HCS structurepeiThreedimensionalWavyHeliospheric2012 . By incorporating the LIS obtained from the GALPROP package moskalenkoProductionPropagationCosmicRay1998 ; strongPropagationCosmicRayNucleons1998a , we demonstrate that this model has potential to provide a unified description of modulation for all four elementary cosmic-ray species, irrespective of charge or mass, and is fitted to Voyager and AMS-02 measurements. Such a unified framework can be further applied to study potential anomalous signals obscured in the GeV-range cosmic-ray spectra.
@@ III.2 Diffusion Model
The three-dimensional Parker’s equation, with the complex structure described above, is solved using a backward-in-time SDE method. In this approach, pseudo-particles are traced backward from the Earth to the heliopause. The evolution of a particle’s position is governed by the SDE yamadaStochasticViewSolar1998 ; zhangMarkovStochasticProcess1999 ; koppStochasticDifferentialEquation2012 ; kapplSOLARPROPChargesignDependent2016 :
@@ IV.1 Machine Learning Model
Figure 2: Deep learning architecture for spectral data processing in space physics, featuring propagation matrix (PropMat) generation and analysis across spatial domains (Galaxy and Heliosphere) using an encoder-decoder framework. The GConv2d operator depicted in the figure denotes a two-dimensional grouped convolution layer.
@@
A matrix generation via controlled multi-scale block operations is used to replace the conventional MLP decoder in PropMat. The module employs PixelShuffle operations to enhance localized correlations across different spatial hierarchies. In this framework, each matrix is dynamically split into independent groups at different hierarchical scales, with elements appear simultaneously in both higher- and lower-level groups. The fully‑connected operations are then applied separately for all groups. This process amplifies interactions among spatially adjacent positions, as they are linked through multiple connections across scales. Therefore, it inherently encourages spatial smoothness in the generated matrices while preserving computational efficiency through sparse group‑wise operations.
@@
$\text{Loss}_{\text{G}}=\sum_{j}\Bigl(\log\frac{\sum_{i}M_{\text{pred}}^{ij}\,\psi_{\mathrm{I}}^{i}}{\psi_{\mathrm{O}}^{j}}\Bigr)^{2}+\sum_{ij}\lambda\log\bigl(M_{\text{dev}}^{ij}\bigr)^{2},$ (17)
where the MSE is evaluated on a logarithmic scale because the spectral values span several orders of magnitude. The hyperparameter $\lambda$, which controls the regularization strength, is assigned with a relatively large value at first and decreased gradually.
@@
where the MSE is also evaluated on a logarithmic scale as in Equation (17). Since the typical energy change during heliospheric modulation is on the order of the modulation potential (approximately 1GeV1\ $\mathrm${GeV} for $|Z|=1$ particles), the energy-bin width in the transformation matrix must be considerably smaller than this scale. We therefore adopt a linear binning scheme with widths of 0.1GeV0.1\ $\mathrm${GeV} or 0.05GeV0.05\ $\mathrm${GeV}.
@@ V Results
Figure 6: Best-fit proton (Top) and antiproton (Bottom) fluxes. The black dash-dotted line corresponds to the LIS, while colored solid and dashed lines display spectra modulated using the FFA and HELPROP, respectively. Line colors match respective experimental measurements, with FFA modulation potentials $\phi$ indicated in the legend.
@@
Figure 6 shows the best-fit local interstellar spectra for protons and antiprotons, along with the modulated spectra for several representative Bartels rotations, compared with Voyager and AMS-02 data. For comparison, the modulation result obtained using the FFA is plotted with dashed lines. The proton fluxes during Bartels rotations 2543, 2554, and 2557 exhibit close agreement, while the antiproton fluxes in these same rotations show significant discrepancies. This flux disparity is primarily attributed to variations in the tilt angle $\alpha$. During the studied period under positive magnetic polarity, as mentioned in Section III, protons reach Earth primarily through the polar region and are thus insensitive to changes in the tilt angle. In contrast, antiprotons detected at Earth are transported mainly via HCS drift. A larger tilt angle extends the drift path, enhances particle losses during transport, and consequently lowers the observed flux. Compared to the FFA, which requires distinct modulation potentials for protons and antiprotons to account for such differences, our model incorporates tilt angle information and successfully reproduces the observed variations for both both particle species.
This unified description of charge-dependent modulation can be further tested through the correlation between antiproton and proton fluxes over the full observation period, as shown in Fig. 7. Such charge-dependent modulation behavior is not limited to a few individual rotations but constitutes a systematic feature throughout the entire solar quiet period. Figure 7 presents the correlation between antiproton and proton fluxes at a kinetic energy of 1 GeV. To mitigate the effects of short-term fluctuations, the data points are shown as 13-Bartels-rotation moving averages, normalized by their respective time-averaged flux values. Our model consistently describes the flux variations of both charge signs across the entire observation interval, and reproduces the correlation between protons and antiprotons. This further proves that the HELPROP model, based on a realistic three-dimensional heliospheric current sheet drift, can quantitatively and uniformly characterize charge-dependent modulation effects, outperforming the force-field approximation that requires independent modulation potentials for different charge signs.
Figure 7: Correlation between antiproton and proton fluxes in the rigidity range of 1.00–2.97 GV, along with the corresponding best-fit result. The data points represent the 13-Bartels-rotation moving averages of observed fluxes, normalized to their respective time-averaged values. Colored points correspond to quiet periods and are included in the fitting, whereas black points represent active periods and are excluded from the fit. The solid black line shows the fitting result.